# Robust $\mathcal{H}_{\infty}$ control and worst-case search in constrained parametric space

### v2

@@ Abstract
Standard $\mathcal{H}_{\infty}$$\mathcal{H}_{2}$ robust control and analysis tools operate on uncertain parameters assumed to vary independently within prescribed bounds. This paper extends their capabilities in the presence of constraints coupling these parameters and restricting the parametric space. Focusing on the worst-case search, we demonstrate – based on the theory of upper-$C^{1}$ functions – the validity of using standard, readily available smooth optimization algorithms to address this nonsmooth constrained optimization problem. Accordingly, we propose to explore the parametric space with either Monte-Carlo sampling or particle swarm optimization, and to subsequently perform local exploitation with Sequential Quadratic Programming to compute Karush-Kuhn-Tucker points. This worst-case search then enables robust controller synthesis: as in the state-of-art algorithm for standard robust control, identified worst-case configurations are iteratively added to an active set on which a non-smooth multi-models optimization of the controller is performed. The methodology is illustrated on a satellite benchmark with flexible appendages, of order 50 with 43 uncertain parameters. We show that the proposed method largely outperforms Monte-Carlo sampling alone, is able to reliably detect even rare worst-case configurations in minutes on a standard laptop, and that the robust controller optimization converges with less than 10 active configurations. Even in the unconstrained case, the proposed framework complements traditional methods, as it scales to plants with many parameters and states and explores the entire parametric space, albeit without formal guarantees of global optimality.
@@ I Introduction
The standard robust structured $\mathcal{H}_{\infty}$ control problem [apkarian2017] is formulated as the min-max optimization problem:
@@
In this paper, we mainly focus on the inner maximization problems, namely search for worst-case performance:
@@
and worst-case stability – stability is implicitly required to ensure finite norms – treated as in [apkarian2015a]:
@@
Classically in the literature, $\mathcal{D}$ is the hypercube $1,1]^{k}$ after normalization of the uncertain parameters, and $T_{zw}(s,K,\delta)$ is modeled under the Linear Fractional Representation (LFR). In this case, MATLAB’s function systune, relying on the non-smooth optimization of [apkarian2015a, apkarian2006], is now widely used in academia and industry to solve problems (1) and (2). More specifically, in [apkarian2015a], problems (3) and (4) are solved with local non-smooth optimization. Indeed, the computation of Clarke subdifferentials [Clarke90] of the spectral abscissa and of the $\mathcal{H}_{\infty}$ norm [Boyd91] enables a bundle method to quickly find worst-case configurations with regard to stability and performance. The control problem (1) is then tackled by combining this worst-case search with the non-smooth controller optimization of [apkarian2006]. This approach relies on the construction of an inner approximation $\mathcal{D}_{a}\subset\mathcal{D}$ composed of so-called active configurations. At each iteration, the controller is tuned so as to be robust to all current active configurations. Then, a worst-case search is performed; if the worst-case configuration does not satisfy the control requirements, it is added to $\mathcal{D}_{a}$ and the algorithm goes back to controller tuning; otherwise, the algorithm can terminate. It is shown that this approach is both more reliable and more efficient than gridding the parametric space, which quickly becomes intractable when the number of parameters increases. In practice, the optimization generally finishes with a small number of active configurations. As this approach relies on local searches, robustness is generally thoroughly verified a posteriori. $\mu$-analysis [doyle1982] was developed for this purpose, generally based on branch-and-bound algorithms [roos2013systems]. It provides guaranteed deterministic lower and upper bounds of the global optimum; a probabilistic variant also exists [Biannic2021, Somers2025]. Monte-Carlo sampling can also guarantee a probability of robustness within predefined thresholds of accuracy and confidence [Tempo1996]. Another common industrial practice is to grid the parametric space with a limited number of samples, typically 100 to 1000 combinations of the parameters bounds. But this approach will miss any worst case occurring between grid points and, unlike Monte-Carlo, does not provide any statistical guarantee.
@@
Our focus is twofold. Firstly, we seek computationally efficient techniques that can be integrated into robust control design, alternating between multi-models controller synthesis and worst-case search as in [apkarian2015a], to solve the standard problem (1) and its multi-objectives formulation (2) in reasonable time. Secondly, we aim to assess their reliability in consistently detecting global worst-case configurations as a validation tool.
@@
The robust control and analysis approaches discussed above do not address the constrained case, at least not in their current implementations. Nevertheless, quantifying robustness can still be approached with Monte-Carlo sampling; samples that do not satisfy the nonlinear constraint are simply discarded. Evaluating the stability or the norm of a linear system at a given parametric configuration is relatively fast, and easily parallelizable; moreover, the number of evaluations required to provide the desired levels of probability and confidence is independent from the number of parameters [Tempo1996]. However, finding a worst case, e.g. to address problems (3) or (4), may require numerous function evaluations and lead to a significant computation time, especially in critical systems for which even rare worst cases must be avoided. This is particularly an issue when the robustness analysis iterates with controller design to perform robust tuning.
@@
In this paper, we demonstrate new properties of the class of upper-$C^{1}$ functions [Spingarn81], and in particular the convergence of Sequential Quadratic Programming (SQP) [Nocedal2006], a standard constrained optimization algorithm, when applied to such functions even though they are not smooth. We deduce that SQP can be applied to the nonsmooth problem (3), and to (4) under an additional assumption. Then, relying on built-in MATLAB functions, we propose to explore the parametric space with either Particle Swarm Optimization (PSO) [Kennedy1995], tackling the nonlinear constraint through a penalty approach [Davenport], or Monte-Carlo sampling (if stochastic guarantees are sought), and subsequently refine the search with SQP to find Karush-Kuhn-Tucker points of the optimization problem. Then, following the approach of [apkarian2015a], this worst-case search method is then combined with multi-models controller optimization [apkarian2006] to perform robust control. To the best of our knowledge, this is the first robust $\mathcal{H}_{\infty}$ control approach that tackles nonlinear constraints on the uncertain parameters.
@@
We emphasize that the proposed techniques, although studied here in the context of the constrained space (5), may also prove valuable in the unconstrained case. Indeed, $\mu$-analysis remains difficult to apply to industrial systems with many uncertain parameters – as illustrated in the recent references [Kassarian2023, Kassarian2024, Martin2023, Martin2025] where systems with 16 to 20 parameters required sensitivity-based model reduction to make the problem tractable, at the expense of model fidelity. In particular, [Sanfedino2025] shows that such $\mu$-analysis can even lead to an overly optimistic worst-case prediction after model reduction. This occurs because the formal robustness guarantees apply only to the reduced model and do not extend to the original system, thus limiting the practical value of sensitivity-based $\mu$-analysis. By contrast, the proposed approach scales well to plants with many states and parameters, and thus can be applied to the full system. Furthermore, the non-smooth optimization of [apkarian2015a], implemented in systune, is local and sometimes misses worst-case configurations, whereas the proposed approach performs a global exploration of the parametric space, albeit without formal guarantees of global optimality.
@@
We illustrate the proposed approach on the robust control of a spacecraft with flexible appendage. In such systems, uncertainties associated with the mechanical properties of flexible appendages must satisfy a condition of the form (5) or otherwise be excluded from robustness analysis (specifically, the residual mass matrix must remain positive definite). While structured $\mathcal{H}_{\infty}$ control has become industrial standard practice among satellite integrators for missions with demanding control requirements [Martin2023, bepicolombo2014, Falcoz2015, Charbonnel2010], this constraint, and more generally the robustness analysis, is generally handled using Monte-Carlo or gridding approaches, which suffer from the aforementioned drawbacks. The proposed methodology offers a practical and scalable alternative that can significantly reduce engineering effort and computational cost.
@@ II-A Properties of upper-$C^{1}$ functions
Definition 1 (Lower- and upper-$C^{1}$ functions [Spingarn81])
@@
[apkarian2015a, Sections III.A., II.B., V.A.] shows that minimization of such upper-$C^{1}$ functions, even when not differentiable everywhere, exhibits similar behavior to smooth functions. In particular, the authors show the convergence of their bundle method, in the search of the worst-case $\mathcal{H}_{\infty}$ norm, even when the cutting plane step is reduced to a standard linesearch (as in smooth optimization). This work follows [Dao2015] that treats of the optimization of lower- and upper-$C^{1}$ functions. Similar considerations are also thoroughly discussed by the same authors in [Apkarian2017wc]. Similarly, in [Oliveira2023], the Frank-Wolfe algorithm, originally designed for smooth functions, is also proven to converge to Clarke-stationnary points when applied to upper-$C^{1}$ functions. In the same spirit, we now show that the upper-$C^{1}$ functions, although not differentiable everywhere, maintain some favorable properties for their minimization. Then, we will demonstrate the convergence of one particular smooth optimization algorithm when applied to upper-$C^{1}$ functions.
@@
a) [Oliveira2023, Proposition 2 and equation (6)] shows that the directional derivative of an upper-$C^{1}$ function $f$ is:
@@
Proposition 1.a) shows that when a smooth optimization algorithm is provided with any subgradient as if it were the gradient of a smooth function, progress can still be achieved at points of nonsmoothness, owing to the regularity property of upper-$C^{1}$ functions. Naturally, the selected direction is not necessarily the steepest descent. Proposition 1.b) ensures that an optimization algorithm can rely on the same optimality conditions as in the smooth case when picking any subgradient (we recall that, for the general case of nonsmooth functions, the optimality condition is $0\in\partial f(x)$, which is a weaker condition).
@@
where $H$ is a symmetric definite positive matrix and $g_{x}\in\partial f(x)$. The quadratic problem $Q(x,H,g_{x})$ can be seen as a local approximation of the Problem (8) where the matrix $H$ may contain information about the curvature of both objective and constraint functions, usually serving as an approximation of the Lagrangian. $Q(x,H,g_{x})$ is originally introduced in methods such as [Han1997] where $f$ is differentiable and $\nabla f(x)$ replaces $g_{x}$.
Finally, for $r\geq 0$, we introduce the merit function $\theta_{r}$ as in [Han1997]:
@@
a) Descent direction. Let $(p,u)$ be a KKT pair of $Q(x,H,g_{x})$. If $p||\neq 0$ and $u||_{\infty}\leq r$, then $p$ is a descent direction for $\theta_{r}(x)$, i.e. $\theta^{\prime}_{r}(x,p)<0$ where $\theta^{\prime}_{r}(x,d)$ denotes the the directional derivative at $x$ in direction d.
@@
a) Cf. Appendix. The proof is a mere adaptation of [Han1997, Theorem 3.1] to the upper-$C^{1}$ case.
@@
As shown in [Clarke90, Oliveira2023], the subdifferential of an upper-$C^{1}$ function $f$ at $x$ is the convex hull formed by the gradients of the active functions $F(\cdot,s)$:
@@
Proposition 1.b) is similar to [Oliveira2023, Proposition 4], where it is shown that ”directional-stationarity”, i.e. the assumption that $f^{\prime}(x,d)\geq 0$ for any $d\in\mathbb{R}^{k}$, implies that all subgradients are 0. Note that directional stationarity does not imply that $x$ minimizes $f(x)$; for this, more assumptions would be required, e.g. convexity of the $F(\cdot,s)$ [Oliveira2023, Theorem 10]. Proposition 2.a) is a generalization of [Han1997, Theorem 3.1], originally for differentiable functions, to upper-$C^{1}$ functions. To the best of our knowledge, Proposition 2.b) is novel, and can also be seen as a specific case of the usual KKT conditions for upper-$C^{1}$ functions. It allows optimality to be checked with any subgradient, in contrast to the general nonsmooth case, where only the existence of one subgradient satisfying KKT conditions is required.
@@ II-B Convergence of sequential quadratic programming (SQP) for upper-$C^{1}$ functions
We now introduce the Sequential Quadratic Programming (SQP) algorithm to solve Problem (8). It is described in Algorithm 1, as in [Han1997] except that we replaced the gradient $\nabla f(x)$ by any subgradient $g_{x}\in\partial f(x)$ in the quadratic problem $Q(x,H,g_{x})$ (equation (10)). See also [Powell1983] for a discussion when the penalty parameter $r$ is variable. We do not detail how the quadratic problem in Step 1 or the step size $\lambda_{k}$ in Step 2 may be solved, or how $H_{k}$ is updated in Step 3, as it does not affect the convergence proof (as long as the conditions described in Algorithm 1 are satisfied)
Algorithm 1 Sequential Quadratic Programming (SQP) algorithm [Han1997] for upper-$C^{1}$ functions
@@
Proposition 3 extends the convergence result of SQP established in [Han1997, Theorem 3.2] for differentiable objective functions to the case of upper-$C^{1}$ objective functions. Besides the upper-$C^{1}$ condition of $f$, the same assumptions as in [Han1997, Theorem 3.2] are considered.
@@
Cf. Appendix. The proof follows that of [Han1997, Theorem 3.2] and adapts it to the case where f is not differentiable but only upper-$C^{1}$. In particular, the upper semi-continuity of the Clarke subdifferential allows to maintain the argument linked to the continuity of the gradient in the original proof. The existence of an accumulation point of the solutions $p_{k}$ of the quadratic problems $Q(x_{k},H_{k},g_{x_{k}})$ also follows from the submonotonicity of the subdifferential of lower-$C^{1}$ functions [Spingarn81] (cf. Lemma 1 in Appendix). ∎
Proposition 2.b) establishes that local minima are necessarily KKT points for any subgradient, and Proposition 3 proves the convergence to KKT points when using any subgradient. Consequently, the standard SQP algorithm can be applied to upper-$C^{1}$ functions, by providing any subgradient where the gradient is expected.
@@
The purpose of our arguments for using smooth optimization is not to claim superiority over specialized nonsmooth optimization methods. Instead, we emphasize that readily available algorithms (e.g., in MATLAB or Python toolboxes) can be applied successfully and produce satisfactory results.

### v4

@@ Abstract
Standard $\mathcal{H}_{\infty}$$\mathcal{H}_{2}$ robust control and analysis tools operate on uncertain parameters assumed to vary independently within prescribed bounds. This paper extends their capabilities in the presence of constraints coupling these parameters and restricting the parametric space. Focusing on the worst-case search, we demonstrate – based on the theory of upper-$C^{1}$ functions – the validity of using standard, readily available smooth optimization algorithms to address this nonsmooth constrained optimization problem. In particular, we prove that the sequential quadratic programming algorithm converges to Karush-Kuhn-Tucker points, and that such conditions are satisfied by any subgradient at a local minimum. This worst-case search then enables robust controller synthesis: identified worst-case configurations are iteratively added to an active set on which a non-smooth multi-models optimization of the controller is performed. The methodology is illustrated on a satellite benchmark with flexible appendages, of order 50 with 43 uncertain parameters. From a practical point of view, we combine the local exploitation proposed above with a global exploration using either Monte-Carlo sampling or particle swarm optimization. We show that the proposed constrained optimization effectively complements Monte-Carlo sampling by enabling fast detection of rare worst-case configurations, and that the robust controller optimization converges with less than 10 active configurations.
@@ I Introduction
The standard robust structured $\mathcal{H}_{\infty}$ control problem [10] is formulated as the min-max optimization problem:
@@
In this paper, we mainly focus on the inner maximization problems, which will then allow us to solve Problems (1) and (2). Namely, search for worst-case performance:
@@
and worst-case stability – stability is implicitly required to ensure finite norms – treated as in [3]:
@@
Classically in the literature, $\mathcal{D}$ is the hypercube $1,1]^{k}$ after normalization of the uncertain parameters, and $T_{zw}(s,K,\delta)$ is modeled under the Linear Fractional Representation (LFR). In this case, MATLAB’s function systune, relying on the non-smooth optimization of [3, 8], is now widely used in academia and industry to solve problems (1) and (2). More specifically, in [3], problems (3) and (4) are solved with local non-smooth optimization. Indeed, the computation of Clarke subdifferentials [22] of the spectral abscissa and of the $\mathcal{H}_{\infty}$ norm [16] enables a bundle method to quickly find worst-case configurations with regard to stability and performance. The control problem (1) is then tackled by combining this worst-case search with the non-smooth controller optimization of [8]. This approach relies on the construction of an inner approximation $\mathcal{D}_{a}\subset\mathcal{D}$ composed of so-called active configurations. At each iteration, the controller is tuned so as to be robust to all current active configurations. Then, a worst-case search is performed; if the worst-case configuration does not satisfy the control requirements, it is added to $\mathcal{D}_{a}$ and the algorithm goes back to controller tuning; otherwise, the algorithm can terminate. It is shown that this approach is both more reliable and more efficient than gridding the parametric space, which quickly becomes intractable when the number of parameters increases. In practice, the optimization generally finishes with a small number of active configurations. As this approach relies on local searches, robustness is generally thoroughly verified a posteriori. $\mu$-analysis [29] was developed for this purpose, generally based on branch-and-bound algorithms [47]. It provides guaranteed deterministic lower and upper bounds of the global optimum; a probabilistic variant also exists [13, 50]. Monte-Carlo sampling can also guarantee a probability of robustness within predefined thresholds of accuracy and confidence [53]. Another common industrial practice is to grid the parametric space with a limited number of samples, typically 100 to 1000 combinations of the parameters bounds. But this approach will miss any worst case occurring between grid points and, unlike Monte-Carlo, does not provide any statistical guarantee.
@@
Our main focus is to seek computationally efficient techniques that can be integrated into robust control design, alternating between multi-models controller synthesis and worst-case search as in [3], to solve the standard problem (1) and its multi-objectives formulation (2) in reasonable time.
@@
The robust control and analysis approaches discussed above do not address the constrained case, at least not in their current implementations. Nevertheless, quantifying robustness can still be approached with Monte-Carlo sampling; samples that do not satisfy the nonlinear constraint are simply discarded. Evaluating the stability or the norm of a linear system at a given parametric configuration is relatively fast, and easily parallelizable; moreover, the number of evaluations required to provide the desired levels of probability and confidence is independent from the number of parameters [53]. However, finding a worst case, e.g. to address problems (3) or (4), may require numerous function evaluations and lead to a significant computation time, especially in critical systems for which even rare worst cases must be avoided. This is particularly an issue when the robustness analysis iterates with controller design to perform robust tuning.
@@
In this paper, we demonstrate new properties of the class of upper-$C^{1}$ functions [51], and in particular the convergence of Sequential Quadratic Programming (SQP) [38], a standard constrained optimization algorithm, when applied to such functions even though they are not smooth. We deduce that SQP can be applied to the nonsmooth problem (3), and to (4) under an additional assumption, to find their Karush-Kuhn-Tucker points. From a practical point of view, we also study the use of Particle Swarm Optimization (PSO) [34], tackling the nonlinear constraint through a penalty approach [27], or Monte-Carlo sampling (if stochastic guarantees are sought) to perform a global exploration of the parameter space, with SQP subsequently ensuring efficient local exploitation. Then, following the approach of [3], this worst-case search method is then combined with multi-models controller optimization [8] to perform robust control. To the best of our knowledge, this is the first robust $\mathcal{H}_{\infty}$ control approach that tackles nonlinear constraints on the uncertain parameters.
@@
We emphasize that the proposed techniques, although studied here in the context of the constrained space (5), may also prove valuable in the unconstrained case. Indeed, $\mu$-analysis remains difficult to apply to industrial systems with many uncertain parameters – as illustrated in the recent references [33, 32, 36, 37] where systems with 16 to 20 parameters required sensitivity-based model reduction to make the problem tractable, at the expense of model fidelity. In particular, [49] shows that such $\mu$-analysis can even lead to an overly optimistic worst-case prediction after model reduction. This occurs because the formal robustness guarantees apply only to the reduced model and do not extend to the original system, thus limiting the practical value of sensitivity-based $\mu$-analysis. By contrast, the proposed approach scales well to plants with many states and parameters (as a gradient-based method), and thus can be applied to the full system. Furthermore, the non-smooth optimization of [3], implemented in systune, is local and sometimes misses worst-case configurations, whereas the proposed approach performs a global exploration of the parametric space, albeit without formal guarantees of global optimality.
@@
We illustrate the proposed approach on the robust control of a spacecraft with flexible appendage. In such systems, uncertainties associated with the mechanical properties of flexible appendages must satisfy a condition of the form (5) or otherwise be excluded from robustness analysis (specifically, the residual mass matrix must remain positive definite). While structured $\mathcal{H}_{\infty}$ control has become industrial standard practice among satellite integrators for missions with demanding control requirements [36, 52, 30, 20], this constraint, and more generally the robustness analysis, is generally handled using Monte-Carlo or gridding approaches, which suffer from the aforementioned drawbacks. The proposed methodology offers a practical and scalable alternative that can significantly reduce engineering effort and computational cost.
@@ II-A Properties of upper-$C^{1}$ functions
Definition 1 (Lower- and upper-$C^{1}$ functions [51])
@@
[3, Sections III.A., II.B., V.A.] shows that minimization of such upper-$C^{1}$ functions, even when not differentiable everywhere, exhibits similar behavior to smooth functions. In particular, the authors show the convergence of their bundle method, in the search of the worst-case $\mathcal{H}_{\infty}$ norm, even when the cutting plane step is reduced to a standard linesearch (as in smooth optimization). This work follows [26] that treats of the optimization of lower- and upper-$C^{1}$ functions. Similar considerations are also thoroughly discussed by the same authors in [11]. Similarly, in [28], the Frank-Wolfe algorithm, originally designed for smooth functions, is also proven to converge to Clarke-stationnary points when applied to upper-$C^{1}$ functions. In the same spirit, we now show that the upper-$C^{1}$ functions, although not differentiable everywhere, maintain some favorable properties for their minimization. Then, we will demonstrate the convergence of one particular smooth optimization algorithm when applied to upper-$C^{1}$ functions.
@@
The directional derivative at $x$ in direction $d$ of $f$ defined as
$f^{\prime}(x,d)=\lim_{t\to 0^{+}}\frac{f(x+td)-f(x)}{t}$
always exists for an upper-$C^{1}$ function [46].
@@
a) [28, Proposition 2 and equation (6)] shows that the directional derivative of an upper-$C^{1}$ function $f$ is:
@@
Proposition 1.a) shows that when a smooth optimization algorithm is provided with any subgradient as if it were the gradient of a smooth function, progress can still be achieved at points of nonsmoothness, owing to the regularity property of upper-$C^{1}$ functions. Naturally, the selected direction is not necessarily the steepest descent. Proposition 1.b) ensures that an optimization algorithm can rely on the same optimality conditions as in the smooth case when picking any subgradient (we recall that, for the general case of nonsmooth functions, the optimality condition is $0\in\partial f(x)$ which is a weaker condition, also called Clarke-stationarity, and requires a specific treatment in nonsmooth optimization algorithms).
@@
where $H$ is a symmetric definite positive matrix and $g_{x}\in\partial f(x)$. The quadratic problem $Q(x,H,g_{x})$ can be seen as a local approximation of the Problem (8) where the matrix $H$ may contain information about the curvature of both objective and constraint functions, usually serving as an approximation of the Lagrangian. $Q(x,H,g_{x})$ is originally introduced in methods such as [31] where $f$ is differentiable and $\nabla f(x)$ replaces $g_{x}$.
Finally, for $r\geq 0$, we introduce the merit function $\theta_{r}$ as in [31]:
@@
a) Descent direction. Let $(p,u)$ be a KKT pair of $Q(x,H,g_{x})$. If $p||\neq 0$ and $u||_{\infty}\leq r$, then $p$ is a descent direction for $\theta_{r}(x)$, i.e. $\theta^{\prime}_{r}(x,p)<0$.
@@
a) Cf. Appendix. The proof is a mere adaptation of [31, Theorem 3.1] to the upper-$C^{1}$ case.
@@
As shown in [22, 28], the subdifferential of an upper-$C^{1}$ function $f$ at $x$ is the convex hull formed by the gradients of the active functions $F(\cdot,s)$:
@@
Proposition 1.b) is similar to [28, Proposition 4], where it is shown that ”directional-stationarity”, i.e. the assumption that $f^{\prime}(x,d)\geq 0$ for any $d\in\mathbb{R}^{k}$, implies that all subgradients are 0. Note that directional stationarity does not imply that $x$ minimizes $f(x)$; for this, more assumptions would be required, e.g. convexity of the $F(\cdot,s)$ [28, Theorem 10]. Proposition 2.a) is a generalization of [31, Theorem 3.1], originally for differentiable functions, to upper-$C^{1}$ functions. To the best of our knowledge, Proposition 2.b) is novel, and can also be seen as a specific case of the usual KKT conditions for upper-$C^{1}$ functions. It allows optimality to be checked with any subgradient, in contrast to the general nonsmooth case, where only the existence of one subgradient satisfying KKT conditions is required.
@@ II-B Convergence of sequential quadratic programming (SQP) for upper-$C^{1}$ functions
We now introduce the Sequential Quadratic Programming (SQP) algorithm to solve Problem (8). It is described in Algorithm 1, as in [31] except that we replaced the gradient $\nabla f(x)$ by any subgradient $g_{x}\in\partial f(x)$ in the quadratic problem $Q(x,H,g_{x})$ (equation (10)). See also [44] for a discussion when the penalty parameter $r$ is variable. We do not detail how the quadratic problem in Step 1 or the step size $\lambda_{k}$ in Step 2 may be solved, or how $H_{k}$ is updated in Step 3, as it does not affect the convergence proof (as long as the conditions described in Algorithm 1 are satisfied)
Algorithm 1 Sequential Quadratic Programming (SQP) algorithm [31] for upper-$C^{1}$ functions
@@
Proposition 3 extends the convergence result of SQP established in [31, Theorem 3.2] for differentiable objective functions to the case of upper-$C^{1}$ objective functions. Besides the upper-$C^{1}$ condition of $f$, the same assumptions as in [31, Theorem 3.2] are considered.
@@
Cf. Appendix. The proof follows that of [31, Theorem 3.2] and adapts it to the case where f is not differentiable but only upper-$C^{1}$. In particular, the upper semi-continuity of the Clarke subdifferential allows to maintain the argument linked to the continuity of the gradient in the original proof. The existence of an accumulation point of the solutions $p_{k}$ of the quadratic problems $Q(x_{k},H_{k},g_{x_{k}})$ also follows from the submonotonicity of the subdifferential of lower-$C^{1}$ functions [51] (cf. Lemma 1 in Appendix). ∎
Proposition 2.b) establishes that local minima are necessarily KKT points for any subgradient, and Proposition 3 proves the convergence to KKT points when using any subgradient. Consequently, the standard SQP algorithm can be applied to upper-$C^{1}$ functions, by providing any subgradient where the gradient is expected. At a local minimum, any subgradient will be zero and thus enable detection of the optimality by the optimization algorithm. Conversely, the SQP algorithm converges to a KKT point, meaning that at least one subgradient is zero; note that convergence to points where all subgradients are zero, is not guaranteed.
@@
The purpose of our arguments for using smooth optimization is not to claim superiority over specialized nonsmooth optimization methods (for example [26, 25] for lower- or upper-$C^{1}$ functions specifically). Instead, we emphasize that readily available algorithms (e.g., in MATLAB or Python toolboxes) can be applied successfully and produce satisfactory results.